Another test

Sergey Lesenko

May 6, 1997

1 Introduction

(PDF), described in Bienz and Cohn [1993Deutsch [1996], and also available from TEX ⟶ DVIPSRokicki [1994] GhostScriptDeutsch [1996] TeX2PDFSojka and Thanh [1996] Haralambous and Rahtz [1995]. We may compare this route with that

1993, and also available from TEX ⟶ DVIPS1994GhostScript1996TeX2PDF1996 1995. We may compare this route with that

Bienz and Cohn, and also available from TEX ⟶ DVIPSRokickiGhostScriptDeutsch TeX2PDFSojka and Thanh Haralambous and Rahtz. We may compare this route with that

Bienz and Cohn, and also available from TEX ⟶ DVIPSRokickiGhostScriptDeutsch TeX2PDFSojka and Thanh Haralambous and Rahtz. We may compare this route with that

Bienz and Cohn [1993], and also available from TEX ⟶ DVIPSRokicki [1994] GhostScriptDeutsch [1996] TeX2PDFSojka and Thanh [1996] Haralambous and Rahtz [1995]. We may compare this route with that

Bienz and Cohn 1993, and also available from TEX ⟶ DVIPSRokicki 1994GhostScriptDeutsch 1996TeX2PDFSojka and Thanh 1996Haralambous and Rahtz 1995. We may compare this route with that

Bienz and Cohn [1993], and also available from TEX ⟶ DVIPSRokicki [1994] GhostScriptDeutsch [1996] TeX2PDFSojka and Thanh [1996] Haralambous and Rahtz [1995]. We may compare this route with that

References

Bienz, T., and R. Cohn, Portable Document Format Reference Manual, Addison Wesley, Reading, MA, 1993.

Deutsch, L., Aladdin ghostscript, 1996, version 4.01, available at

  ftp://ftp.cs.wisc.edu/pub/ghost/aladdin  .

Haralambous, Y., and S. Rahtz, L^{A}TEX, hypertext and PDF, or the entry of TEX into the world of hypertext, TUGboat, 16(2), 162–173, 1995.

Rokicki, T., Dvips: A TEX driver, 1994, version 5.58, available at

  ftp://labrea.stanford.edu/pub/tex  .

Sojka, P., and H. Thanh, The JOY of TeX2PDF — acrobatics with an alternative to DVI format, TUGboat, 17(3), 1996.